OpenStudy (anonymous):

Solve for h: h + 5 = h – 3

terenzreignz (terenzreignz):

Right, so, by virtue of our addition property of equality, we can add -h to both sides and preserve equality: h + 5 + (-h) = h - 3 + (-h), and this yields: 5 = -3, which simply cannot be. Therefore, there is no solution, there is no value for h that would make this equation hold.

OpenStudy (anonymous):

ohhh thanks:)